ABC is a right triangle whose hypotenuse AC is 8 inches long. Given that the area of ABC is 13.9 square inches, find the area of similar DEF whose hypotenuse DF is 20 inches long.
Answer by edjones(8007)

.5ab=13.9
a=27.8/b
a^2+b^2=8^2
(27.8/b)^2+b^2=64
(772.82/b^2)+b^2=64
772.82+b^4=64b^2
b^4-64b^2+772.82=0
b=4.02, a=6.92 (or vice versa) TI-89 calculator
4.02* 20/8 = 10.05
6.92* 20/8 = 17.3
.5*10.05*17.3=86.9 in^2
.
Ed